OpenStudy (anonymous):

what is the radius and the center of the equation? x2+y2+14x-12y+4=0

hero (hero):

Complete the square

OpenStudy (anonymous):

how do i complete the square @Hero

OpenStudy (anonymous):

This is circle's equation . We will compare it with x^2+y^2+2gx+2fy+c=0 then We get g=7 , f=-6 , c=4 To know the centre of this circle We will write (-g,-f) so We get (-7,6) as the centre of this circle n Radius will be calculated by this formula \[\sqrt{{g}^2+{f}^2-c}\] so The radius =\[\sqrt{6^2+(-7)^2+4}=\sqrt{89}=9.43 \] Answer 100 %

OpenStudy (anonymous):

thank u @Shashank4488

OpenStudy (anonymous):

Check Answer in Book ?

hero (hero):

That's not how he's expected to do it @Shashank4488

OpenStudy (anonymous):

1000 % Confirmed Answer